\begin{thebibliography}{1}
\bibitem{jw2020} K. Jaber and F. O. Wagner, Largeness and equational probability in groups, Ann. Math. Blaise Pascal, 27 (2020) no. 1, 1--17.
\bibitem{jw2000} K. Jaber and F. O. Wagner,  Largeur et nilpotence, Comm. Algebra, 28  (2000) no. 6, 2869--2885.
\bibitem{Kh} E. I. Khukhro, Nilpotent groups and their automorphisms. De Gruyter Expositions in Mathematics, 8. Walter de Gruyter \& Co., Berlin, 1993.
\bibitem{LS} M. Larsen and A. Shalev, Words, Hausdorff dimension and randomly free groups, Math. Ann. 371 (2018), no. 3-4, 1409--1427.
\bibitem{LP} L. L\'evai and L. Pyber,  Profinite groups with many commuting pairs or involutions,  Arch. Math. (Basel) 75 no. 7 (2000) 1--7.
\bibitem{MK} N. Yu. Makarenko and E. I. Khukhro,  Groups with largely splitting automorphisms of orders three and four, Algebra and Logic 42 (2003), no. 3, 165–-176.
\bibitem{Ma} W. A. Manning, Groups in which a large number of operators may correspond to their inverses, Trans. Amer. Math. Soc. 7 (1906) no. 2, 233--240.
\bibitem{SAE} M. Soleimani Malekan, A. Abdollahi and M. Ebrahimi, Compact groups with many elements of bounded order, J. Group Theory, 23 (2020) no. 6 991--998.
\bibitem{R}  D. J. S. Robinson, Finiteness conditions and generalized soluble groups, Part 2, Ergebnisse der Mathematik und ihrer Grenzgebiete, Band 63. Springer-Verlag, New York-Berlin, 1972. 
\bibitem{Ko} Unsolved problems in group theory. The Kourovka notebook. 19th edition. Edited by V. D. Mazurov and E. I. Khukhro. Russian Academy of Sciences Siberian Division, Institute of Mathematics, Novosibirsk, 2019. 
\end{thebibliography}

\documentclass{amsart}
\usepackage{amssymb,hyperref} 
\newtheorem{thm}{Theorem}[section]

\newtheorem{lem}[thm]{Lemma}
\newtheorem{prop}[thm]{Proposition}
\theoremstyle{definition}

\newtheorem{rem}[thm]{Remark}

\newtheorem{qu}[thm]{Question}
\newtheorem{con}[thm]{Conjecture}
\numberwithin{equation}{section}

\newcommand{\set}[1]{\left\{#1\right\}}

\newcommand{\pr}[1]{\left(#1\right)}
\newcommand{\valu}[1]{\left\langle #1\right\rangle}
\newcommand{\ra}[2]{#1\rightarrow #2}
\newcommand{\ma}[2]{#1\mapsto #2}
\begin{document}

\title[Profinite groups with many elements of bounded order]{Profinite groups with many elements of bounded order}%

\author[A. Abdollahi]{Alireza Abdollahi}%
\address{Department of Pure Mathematics, Faculty of Mathematics and Statistics, University of Isfahan, Isfahan 81746-73441, Iran.} 
\email{a.abdollahi@math.ui.ac.ir}%
\author[M. Soleimani Malekan]{Meisam Soleimani Malekan}%
\address{Department of Pure Mathematics, Faculty of Mathematics and Statistics, University of Isfahan, Isfahan 81746-73441, Iran; 
	Institute for Research in Fundamental Sciences, School of Mathematics, Tehran, Iran.} 
\email{msmalekan@gmail.com}

\subjclass[2010]{20E18; 20P05, 43A05}%
\keywords{Profinite groups; elements of bounded order; subsets with positive Haar measures; large subsets}%

\begin{abstract}
L\'evai  and  Pyber proposed the following as a conjecture:\\  
 Let $G$ be a profinite group such that the set of solutions of the equation $x^n=1$
has positive Haar measure. Then $G$ has an open subgroup $H$ and an element $t$ such that all
elements of the coset $tH$ have order dividing $n$ (see Problem 14.53 of [The Kourovka Notebook, No. 19, 2019]). \\
We define a constant $c_n$ for all finite groups and prove that the latter conjecture is equivalent with a conjecture saying $c_n<1$. Using the latter equivalence  we observe that correctness of L\'evai and Pyber conjecture  implies the existence of the universal upper bound $\frac{1}{1-c_n}$ on the index of generalized Hughes-Thompson subgroup $H_n$ of finite groups whenever it is non-trivial. It is known that the latter is widely open even for all primes $n=p\geq 5$.  For odd $n$ we also prove that L\'evai and Pyber conjecture is equivalent to show that  $c_n$ is less than $1$ whenever $c_n$ is only computed on finite solvable groups. \\
The validity of the conjecture has been proved in [Arch. Math. (Basel) 75 (2000) 1-7] for $n=2$. Here we confirm the conjecture for $n=3$. 
\end{abstract}
\maketitle
\section{\bf Introduction and Results}

Let $G$ be a Hausdorff compact group. Then $G$ has a unique normalized Haar measure denoted by ${\mathbf m}_G$. In general, the question of weather the interior of every non-empty measurable subset of $G$ with positive Haar measure is non-empty has negative answer even if $G$ is profinite (see e.g. \cite{LP}). However the same question for subsets defined by words is still open. In \cite{LP} the following conjecture is proposed.   

\begin{con}\label{l-pcon}{\rm (Conjecture 3 of \cite{LP}, Problem 14.53 of \cite{Ko})}   
Let $G$ be a profinite group such that the set $X_n(G)$ of solutions of the equation $x^n=1$ in $G$  has positive Haar measure. Then $G$ has an open subgroup $H$ and an element $t$ such that all elements of the coset $tH$ have order dividing $n$.
\end{con} 

The validity of Conjecture \ref{l-pcon} has been proved in \cite{LP} for $n=2$. In \cite{SAE} it is shown that the conjecture is valid for $n=2$ even if $G$ is   Hausdorff compact.  It is also proved in \cite{SAE} that if $X_3(G)$ has positive Haar measure in a compact group $G$, then $G$ contains an open normal subgroup which is $2$-Engel. Here we confirm Conjecture \ref{l-pcon} for $n=3$.
To do so, we first show that Conjecture \ref{l-pcon} is equivalent to the following one. We need the following notation in the statement of the conjecture. For an arbitrary group $K$ and an automorphism $\phi$ of $K$ of order dividing a positive integer $n$, define 
$$X_{n,\phi}(K):=\left\{x\in K \;|\; xx^\phi x^{\phi^2}\cdots x^{\phi^{n-1}}=1\right\}.$$ 
 The automorphism group of $K$ will be denoted by ${\rm Aut}(K)$.

\begin{con}\label{sup<1}
	  $$ c_n:=\sup \pr{\left\{\frac{|X_{n,\phi}(H)|}{|H|}: H\,
	\text{\rm is a finite group and}\;\; \phi\in{\rm Aut}(H), \; \phi^n={\rm id}\right\} \setminus\set{1}}<1.$$
\end{con}

 It is known that Conjecture \ref{sup<1} is valid for $n=2$ and the supremum $c_2$ is $\frac{3}{4}$ (see \cite{Ma}). We shall prove that $c_3<1$. 

If $n$ is odd, using \cite[Theorem 1.10]{LS}, we prove that Conjecture \ref{l-pcon} is equivalent to the following. Here we denote by $\mathcal{S}$ the class of finite solvable groups. 
\begin{con}\label{supS<1}
	$$ c^{\mathcal{S}}_{2n+1}:=\sup \pr{\left\{\frac{|X_{n,\phi}(H)|}{|H|}: H\in \mathcal{S}\,
		\text{\rm and}\;\; \phi\in{\rm Aut}(H), \; \phi^{2n+1}={\rm id}\right\} \setminus\set{1}}<1.$$
\end{con}

We conclude this section by noting  how above conjectures are difficult. First we note that  
\begin{prop}
	If Conjecture \ref{sup<1} is true for $n$ i.e. $c_n<1$, then for any finite group $G$ with $H_n(G):=\langle x\in G \;|\; x^n\neq 1\rangle\neq 1$, $|G:H_n(G)|<\frac{1}{1-c_n}$.
\end{prop}
\begin{proof}
	Since $H_n(G)\neq 1$, $X_n(G)\neq G$. Now the proof follows from the fact that $G\setminus H_n(G) \subset X_n(G)$ and the definition of $c_n$. 
\end{proof}

\begin{rem}
	The problem of finding a universal bound from above depending only on $p$ for  the index $|G:H_p(G)|$  of the Hughes-Thompson subgroup $H_p(G)$ of finite $p$-groups $G$ with $H_p(G)\neq 1$ is widely open for all $p\geq 5$ (see \cite[Chapter 7]{Kh} for the history and results on the problem).
\end{rem}

We finish this section with the following two questions.

\begin{qu}
	Suppose that $n$ is a positive integer such that there exists a positive integer $k_n$ depending only on $n$ such that $|G:H_n(G)|\leq k_n$ for all finite groups $G$ with $1\neq H_n(G)$. Is it true that $c_n<1$? The same question whenever $n$ is prime.
\end{qu}

\begin{qu}
	Let $n>1$ be a positive integer such that $c_d<1$ for all prime power divisors $d$ of $n$. Is it true that $c_n<1$?
\end{qu}

\section{\bf Profinite groups}

We denote the normalized Haar measure of a compact group $G$ by ${\mathbf m}_G$, and we will simply write ${\mathbf m}$ if there is no ambiguity. \\

The following easy lemma will be used in the sequel without referring to it. 
 
\begin{lem}{\rm (cf. \cite[Lemma 2.5]{jw2020})}\label{n-large}
	Let $G$ be a compact group and $A\subseteq G$ be a measurable subset. Assume that ${\mathbf m}(A)\geq 1-\epsilon$, then ${\mathbf m}\pr{\bigcap_{k=1}^ng_kA}\geq 1-n\epsilon$ for all $g_1,\dotsc,g_n\in G$. The similar result with  strict inequalities holds. 
\end{lem}
\begin{proof}
	By induction on $n$, we prove the result. For $n=1$, it holds as the measure is left-invariant. Assume that the result is true for $n$; therefore 
	\begin{align*}
		{\mathbf m}\pr{\bigcap_{k=1}^{n+1}g_kA}&={\mathbf m}\pr{\bigcap_{k=1}^{n}g_kA}+{\mathbf m}(g_{n+1}A)-{\mathbf m}\pr{\pr{\bigcap_{k=1}^{n}g_kA}\cup(g_{n+1}A)}\\
		&\geq (1-n\epsilon)+(1-\epsilon)-1,\quad\text{by the induction hypothesis,}\\
		&=1-(n+1)\epsilon.
	\end{align*}
\end{proof}

\begin{lem}
	\label{relation} Let $G$ be a compact group and $\phi$ be a continuous automorphism of $G$ of order dividing $n$. Denote by $G\rtimes\valu{\phi}$ the semidirect product of $G$ by $\valu{\phi}$. Then: 
	\begin{itemize}
		\item[(i)] $X_{n,\phi}(G)$ has nonempty interior if and only if $X_n(G\rtimes\valu{\phi})$ has nonempty interior. 
		\item[(ii)] If $X_{n,\phi}(G)$ has positive Haar measure then $X_n(G\rtimes\valu{\phi})$ has positive Haar measure. 
	\end{itemize}
\end{lem}
\begin{proof}
	It follows from the equality $X_n(G\rtimes\valu{\phi})\cap G\phi^{-1}=X_{n,\phi}(G)\phi^{-1}$.
\end{proof}
\begin{prop}\label{cj implies cj}
 Conjecture \ref{l-pcon} implies  Conjecture \ref{sup<1}.
\end{prop}
\begin{proof}
	If $n$ is such that  Conjecture \ref{sup<1} is not valid, then there   exist sequences $(G_k)$ of finite groups and $(\phi_k)\in\prod_{k=1}^\infty\text{Aut}(G_k)$ such that  $\phi_k^n=1$  and 
	\[ 0<\prod_{k=1}^\infty\frac{|X_{n,\phi_k}(G_k)|}{|G_k|}<1.\]
	Consider the cartezian product $G=\prod_kG_k$ which is clrealy profinite. Then $\phi=(\phi_k)$ is an automorphism of  $G$ of order dividing $n$. It is clear that the measure of $X_{n,\phi}(G)$ is equal to $\prod_k\frac{|X_{n,\phi_k}(G_k)|}{|G_k|}$ and its interior is empty, so by Lemma \ref{relation}, $X_n(G\rtimes\valu{\phi})$ has positive Haar measur and empty interior showing that Conjecture \ref{l-pcon} is not valid. 
\end{proof}

The following lemma will be used in the proof that  ``Conjecture \ref{sup<1} implies  Conjecture \ref{l-pcon}". 
We write ``$N\trianglelefteq_o G$" whenever $N$ is a normal and open subgroup of $G$.
\begin{lem}\label{sup}
	Let $A$ be a closed subset of a profinite group with positive Haar measure and $M$ be any normal open subgroup of $G$. If $\mathcal{X}$ is the set of all normal open subgroups of $G$ contained in $M$. Then 
	\begin{align*}
		\sup\left\{\frac{{\mathbf m}(Ng\cap A)}{{\mathbf m}(N)}: g\in G, N\in {\mathcal X} \right\}=1
	\end{align*}
\end{lem}
\begin{proof} 
	Let $N\in {\mathcal X}$ be such that $r:=|G:N|$. If $s$ is the number of cosets of $N$ which intersect $A$, then 
	\begin{align}
		\label{ine1}
		(r-s) {\mathbf m}(N)\leq 1-{\mathbf m}(A).
	\end{align}
	On the other hand, assume that ${\mathbf m}(Nx\cap A)=\max\{{\mathbf m}(Ng\cap A): g\in G\}$, so 
	\begin{align}\label{ine2}
		{\mathbf m}(A)\leq s{\mathbf m}(Nx\cap A)
	\end{align}
It follows from inequalities (\ref{ine1}) and (\ref{ine2}) that  
	\begin{align*}
		\frac{{\mathbf m}(A)}{1-{\mathbf m}(A)}\frac{r-s}{s}\leq\frac{{\mathbf m}(Nx\cap A)}{{\mathbf m}(N)}.
	\end{align*}
	Since $A$ is closed, ${\mathbf m}(A)=\inf\left\{\frac{|AK/K|}{|G/K|}: K\trianglelefteq_o G \right\}$. Now since $\frac{|AK/K|}{|G/K|}\leq \frac{|AN/N|}{|G/N|}$, whenever $K\leq N$ are normal subgroups of $G$ of finite index, it follows that  
		 ${\mathbf m}(A)=\inf\left\{\frac{|AN/N|}{|G/N|}: N\in {\mathcal X} \right\}$. Since $\frac{r}{s}=\frac{|G/N|}{|AN/N|}$, the result now follows from the last inequality. 
\end{proof}
For any positive integer $n$ and any class  $\mathcal{X}$ of finite groups, we denote by $c^{\mathcal{X}}_n$ the following positive real number which is at most $1$:
$$ \sup \pr{\left\{\frac{|X_{n,\phi}(H)|}{|H|}: H\in \mathcal{X} \,
	 \textbf{\rm and}\;\; \phi\in{\rm Aut}(H), \; \phi^n={\rm id}\right\} \setminus\set{1}}.$$
\begin{prop}\label{**}
Assume $c^{\mathcal{X}}_n<1$ and suppose that	$G$ is a profinite group. Let $M$ be a normal open subgroup of $G$ and $\phi$ be a continues automorphism of $M$ such that $N^\phi\subset N$ for all normal open subgroups $N$ of $G$ contained in $M$ and $M/N \in {\mathcal X}$. Then 
	${\mathbf m}_M(X_{n,\phi}(M))\leq c^{\mathcal{X}}_n$ if $X_{n,\phi}(M)\neq M$.
\end{prop}
\begin{proof}
	Seeking a contradiction, let us suppose ${\mathbf m}_M(X_{n,\phi}(M))> c^{\mathcal{X}}_n$. Let $N$ be a normal open subgroup of $G$. Consider the following  automorphism of $M/(M\cap N)$,
	\[\bar\phi:\frac M{M\cap N}\rightarrow\frac M{M\cap N}, \quad \bar x^{\bar\phi}:=\overline{x^\phi}\]
	We have 
	\[{\mathbf m}_M\pr{X_{n,\phi}(M)}\leq \frac{|X_{n,\bar\phi}(M/(M\cap N))|}{|M/(M\cap N)|}\]
	the inequality holds because if $x\in X_{n,\phi}(M)$, then $\bar x\in X_{n,\bar\phi}(M/(M\cap N))$. Since $c^{\mathcal{X}}_n<1$, so $X_{n,\bar\phi}(M/(M\cap N))=M/(M\cap N)$, whence $\prod_{k=0}^{n-1}x^{\phi^k}\in N$ for all $x\in M$. Since $\bigcap\{N\trianglelefteq_o G: N\leq M\}=\{1\}$,  $\prod_{k=0}^{n-1}x^{\phi^k}=1$ for all $x\in M$, i.e. $X_{n,\phi}(M)=M$.
\end{proof}

We are now in a position to prove the following:
\begin{prop}\label{cj <- cj}
Conjecture \ref{sup<1} implies Conjecture \ref{l-pcon}.
\end{prop}
\begin{proof}
	Let $G$ be a profinite group such that $\textbf{m}_G(X_n(G))>0$. By Lemma \ref{sup}, there exist a normal open subgroup $M$ and $g\in X_n(G)$, such that 
	\begin{align}\label{*}
		c_n<\frac{{\mathbf m}_G\pr{Mg \cap X_n(G)}}{{\mathbf m}_G(M)}=\frac{{\mathbf m}_G\pr{M\cap X_n(G)g^{-1}}}{{\mathbf m}_G(M)}.
	\end{align} 
	Put $\phi:\ra MM$, $\ma x {gxg^{-1}}$. Since $g^n=1$, $\phi^n=1$ and  
	\[ M\cap X_n(G)g=X_{n,\phi}(M)\]
	and the inequality \ref{*} means that ${\mathbf m}_M(X_{n,\phi}(M))>c_n$, so by Proposition \ref{**}, $X_{n,\phi}(M)=M$, whence $Mg\subseteq X_n(G)$. 
\end{proof}

\begin{prop} \label{S->lp}
For odd $n$, Conjecture \ref{supS<1} implies Conjecture \ref{l-pcon}.
\end{prop}
\begin{proof}
Let $G$ be a profinite group such that ${\mathbf m}_G(X_n(G))>0$. By \cite[Theorem 1.10]{LS} and Lemma \ref{sup}, there exist a normal open subgroup $M$ and $g\in X_n(G)$ such that 
such that $M/N$ is finite solvable for all open normal subgroups $N$ of $G$ contained in $M$, and
	\begin{align}\label{*S}
	c^{\mathcal{S}}_n<\frac{{\mathbf m}_G\pr{Mg \cap X_n(G)}}{{\mathbf m}_G(M)}=\frac{{\mathbf m}_G\pr{M\cap X_n(G)g^{-1}}}{{\mathbf m}_G(M)}.
	\end{align} 
	Put $\phi:\ra MM$, $\ma x {gxg^{-1}}$. Since $g^n=1$, $\phi^n=1$ and  
	\[ M\cap X_n(G)g=X_{n,\phi}(M)\]
	and the inequality \ref{*S} means that ${\mathbf m}_M(X_{n,\phi}(M))>c^{\mathcal{S}}_n$, so by Proposition \ref{**}, $X_{n,\phi}(M)=M$, whence $Mg\subseteq X_n(G)$. 
\end{proof}

\section{\bf Compact groups with splitting automorphisms of order $3$} \label{large}

In this section we prove $c_3<1$.

\begin{thm}\label{3split}
	Let $G$ be a compact group and $\alpha$ be an automorphism of $G$ such that $\alpha^3={\rm id}$ and the set $X=\{x\in G \;|\; x x^\alpha x^{\alpha^2}=1\}$ is measurable with ${\mathbf m}(X)>\frac{15}{16}$. Then $X=G$.
\end{thm}
\begin{proof}
First we prove that $G$ is $2$-Engel. The proof is similar to an argument used in the proof of \cite[Theorem 4.4]{SAE}. We give the proof for the reader's convenience.
For any $a,b\in G$ we must prove that $[a,b,b]=1$.  Consider the set 
$$M:=X \cap b^{-1} X \cap aX \cap a^{-1}X \cap ab^{-1}X \cap ba^{-1}X \cap abX \cap b^{-1} a^{-1}X.$$
 Since ${\mathbf m}(M)> \frac{1}{2}$, there exists $x\in X$ such that 
 \begin{align*}
 1&=(x\alpha^{-1})^3=(bx\alpha^{-1})^3=(ax\alpha^{-1})^3=(a^{-1}x\alpha^{-1})^3=(ab^{-1}x\alpha^{-1})^3\\
 &=(ba^{-1}x\alpha^{-1})^3=(abx\alpha^{-1})^3= (b^{-1}a^{-1}x\alpha^{-1})^3,
 \end{align*}
 where we are writing in the semidirect product $G\rtimes \langle \alpha \rangle$ by noting that 
 $g\in X$ if and only if $(g\alpha^{-1})^3=1$ in $G\rtimes \langle \alpha \rangle$.  Now \cite[Lemma 4.1]{SAE} implies that $[a,b,b]=1$.\\

Let $X^{-1}=\{x^{-1} \;|\; x\in X\}$ and $Y=X \cap X^{-1}$.  Then ${\mathbf m}(Y)>\frac{7}{8}$. Note that for all $x\in Y$, 
$$\langle x,x^\alpha,x^{\alpha^2}\rangle=\langle x,x^\alpha \rangle=\langle x^\alpha,x^{\alpha^2}\rangle=\langle x,x^{\alpha^2}\rangle \;\; {\rm are \; all \; abelian.} \;\;\;(*)$$ By \cite[Theorem 7.15 (iv)]{R}, $$[a,b,c]=[a,c,b]^{-1} \;\; {\rm for \; all} \; a,b,c\in G. \;\; (**)$$ 
It follows from $(*)$ and $(**)$ that 
$$[a,b,c]=1 \; {\rm for \; all \;} a,b,c\in\{x,x^\alpha,x^{\alpha^2},h\}, \;\; {\rm for \; all\;} h \in G. \;\; (1)$$

Let $g$ be an arbitrary element of $G$. Consider the set $Z=Y \cap g^{-1} Y$. Since ${\mathbf m}(Z)>\frac{3}{4}$, $Z\neq \varnothing $ and for all  $x\in Z$ we have
$$xx^\alpha x^{\alpha^2}=gx (gx)^\alpha (gx)^{\alpha^2}=1. \;\;\;(2)$$ 
Using $(1)$ and $(2)$ and since $G$ is nilpotent of class at most $3$ (see e.g. \cite[Corollary 3 page 45]{R}), it follows that 
$$(gg^\alpha g^{\alpha^2})^{-1}=[x,g^\alpha] [x,g^{\alpha},g^{\alpha^2}] [g^{\alpha^2},x^{\alpha^2}] \;\; {\rm for \; all\;} x \in Z \;\; (3)$$
Now consider $W=Z \cap x_0^{-1}Z$ for some $x_0\in Z$. Since ${\mathbf m}(W)>\frac{1}{2}$, $W$ is nonempty and for all  $y \in W$ we have 
 $$(gg^\alpha g^{\alpha^2})^{-1}=[x_0y,g^\alpha] [x_0y,g^{\alpha},g^{\alpha^2}] [g^{\alpha^2},(x_0y)^{\alpha^2}]=[y,g^\alpha] [y,g^{\alpha},g^{\alpha^2}] [g^{\alpha^2},y^{\alpha^2}].$$
Note that since $G$ is nilpotent of class at most $3$, commutators of weight $3$ are central. It follows from $(3)$ that 
 $$(gg^\alpha g^{\alpha^2})^{-1}=[x_0,g^\alpha,y][g^{\alpha^2},x_0^{\alpha^2},y^{\alpha^2}] \;\; {\rm for \; all\;} y\in W. \;\; (4)$$
Now consider $T=W \cap y_0^{-1} W $ for some $y_0\in W$. Since ${\mathbf m}(T)>0$, there exists $z\in W$ such that $y_0z\in W$. It follows from $(4)$ that 
$$(gg^\alpha g^{\alpha^2})^{-1}=[x_0,g^\alpha,y_0z][g^{\alpha^2},x_0^{\alpha^2},(y_0z)^{\alpha^2}]=[x_0,g^\alpha,y_0][g^{\alpha^2},x_0^{\alpha^2},y_0^{\alpha^2}].$$
Since commutators of weight $3$ are central in $G$, 
$$[x_0,g^\alpha,y_0z][g^{\alpha^2},x_0^{\alpha^2},(y_0z)^{\alpha^2}]=[x_0,g^\alpha,y_0][g^{\alpha^2},x_0^{\alpha^2},y_0^{\alpha^2}] [x_0,g^\alpha,z][g^{\alpha^2},x_0^{\alpha^2},z^{\alpha^2}].$$
Therefore, $$[x_0,g^\alpha,z][g^{\alpha^2},x_0^{\alpha^2},z^{\alpha^2}]=1$$
and so, as $z\in W$, it follows from $(4)$ that $gg^\alpha g^{\alpha^2}=1$. This completes the proof.
\end{proof}

\begin{rem}\label{rm1}
	Using the first part of our proof of Theorem \ref{3split} and applying \cite[Theorem 6.5]{jw2020} (cf \cite[Th\'eor\`eme 5]{jw2000}) the number $\frac{15}{16}$ can be reduced to $\frac{7}{8}$. 
\end{rem}

\begin{rem}
	In \cite{MK} groups $G$ having an automorphism $\phi \in {\rm Aut}(G)$ with $\phi^n=1$ such that $X_{n,\phi}(G)$ is a large
 set in the sense of \cite{jw2000} for $n=3,4$ are studied. For the case $n=3$ it is proved in \cite{MK} that $G=X_{3,\phi}(G)$. For the case $n=4$, it is proved that a normal solvable subgroup $H$ of $G$ of derived length $d\geq 3$ is nilpotent of class at most $\frac{9^{d-2}+1}{2}$. It is interesting to know if the same latter result is valid  we replace ``largness" of $X_{n,\phi}(G)$ by  weaker condition ``$k$-largness" for some $k$ in the sense of \cite{jw2020}. If the latter is valid, then for compact groups $G$ with ${\mathbf m}_G(X_{4,\phi}(G))>1-\frac{1}{k}$ and $\phi$ is continuous   by Lemma \ref{n-large} we have the same results.  
\end{rem}

\begin{thm}\label{c3<7/8}
	$c_3 < 1$.
\end{thm}
\begin{proof}
It follows from Theorem \ref{3split}.	
\end{proof}

We now see that  Conjecture \ref{l-pcon} of  L\'evai  and  Pyber is true for $n=3$.

\begin{thm}
	Let $G$ be a profinite group such that the equation $x^3=1$ holds on a set with positive Haar measure. 
	Then the solutions set of the equation $x^3=1$ has non-empty interior.
\end{thm}
\begin{proof}
It follows from Theorem \ref{c3<7/8} and Proposition \ref{cj <- cj}, 	
\end{proof}

\end{document}